OpenStudy (anonymous):

Part 1: Is the following equation true? (2 points) Part 2: Use complete sentences to explain the properties used in making your decision. (6 points) 3 log 2 x + one–half log 2 y – 3 log 2 z = log 2 x-cubed times square root of y all over z–cubed

OpenStudy (anonymous):

\[3 \log_{2}x + 1/2 \log_{2}y - 3 \log_{2}z=\log_{2} \frac{ x^3\sqrt{y} }{ z^3 }\]

OpenStudy (anonymous):

Medal and fan for correct answer.

OpenStudy (anonymous):

@Coralgirl @phi @Potatoes.ramu @ranga @jdoe0001 please, someone help!

OpenStudy (anonymous):

@eliassaab @Frostbite @agent0smith

OpenStudy (anonymous):

This is the last question on my homework and I'm so confused.

OpenStudy (anonymous):

@Zarkon @phi @RadEn

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

Guys one of our OS pal need help. Please help

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Luigi0210 @thomaster @

OpenStudy (anonymous):

Come on guys! Please help! I have to do my homework from other classes and I can't until this is done!

OpenStudy (anonymous):

No worries :) they will sure come for help!

OpenStudy (anonymous):

@Coralgirl do you know how to do this?

OpenStudy (anonymous):

Unfortunately I don't :( but dont worry someone will soon now start helping you

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@Mr.ClayLordMath

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@Loser66

OpenStudy (compassionate):

@ranga @dmezzullo @tester97 @jdoe0001 @cwrw238 Get in here.

OpenStudy (anonymous):

LOL so many names but nobody is coming!

OpenStudy (anonymous):

Thanks compassionate. See? @jontonbon all is here to help you. Look @Compassionate one of the amazing OS ambassador also calling all for you to help. Dont you at all, they will be here soon

OpenStudy (anonymous):

You said that 20 minutes ago.

OpenStudy (anonymous):

Dont you worry at all! they will be here soon. They are all must be stuck. But some of them already like Compassionate, agentsmith and tester

OpenStudy (anonymous):

Yes, but they aren't helping, so they are of no use.

OpenStudy (anonymous):

But they are all trying help you.

OpenStudy (tester97):

This is what i got when i googled your question im not sure if this is what your are looking for but here 3log(base2)x + ½log(base2)y - 3log(base2)z use the power property of logs first: log(base2)x³ + log(base2)√y - log(base2)z³ again, use properties of logs where addition means multiply and subtraction means divide: log(base2) [(x³√y)/z³] that's it! ;)

OpenStudy (anonymous):

It's better than nothing, so thank you.

OpenStudy (tester97):

Sometimes i google my questions before posting them to see if they have already been answered.

OpenStudy (tester97):

And you're welcome ^_^

OpenStudy (anonymous):

I did google it but it was coming up with a bunch of random junk, maybe I didn't search hard enough.

OpenStudy (anonymous):

But is the equation true? @tester97

OpenStudy (agent0smith):

Sorry, I wasn't paying attention. The equation is true, i can tell by looking at it, you can use log laws to prove it.

OpenStudy (tester97):

well @jontonbon if |@agent0smith said its true then it must be xD